OpenStudy (danikasosa):

If f(x) = 5x – 3 and g(x) = 3x – 3, find f(x) – g(x). A. 2x B. 8x – 6 C. 2x – 6 D. 8x

OpenStudy (danikasosa):

Somone please help me lmao

OpenStudy (yournightmare):

C. 5-3= 2 and -3-(-3) = 6

OpenStudy (yournightmare):

does that help?

OpenStudy (will.h):

(5x - 3) - (3x -3) I would suggest you to distribute the negative 1. So therefore We would have 5x -3 -3x + 3 Now Combine alike terms 5x -3x = 2x -3 + 3 = 0 So the final result is 2x

OpenStudy (yournightmare):

no its 6 because two negatives make a positive. taking -3 and subtracting 3 would be 6. its not adding three it's subtracting.

OpenStudy (will.h):

Well -3 -(-3) is 0

OpenStudy (yournightmare):

yeah i didnt mean to put that XD

OpenStudy (yournightmare):

im a little tired and stressed. But its -3-3 which is 6